OpenStudy (anonymous):

What's wrong? 1+2+4+8+16+..... =? if 2* (1+2+4+8+16+....) -1 (1+2+4=8+16+...) = (1+2+4+8+16+...) However, if we distribute 2 into terms on the first term, we have 2+4+8+16+..... so that the first term is 1 less than the second term. And the expression =-1. How?? Please, explain

ganeshie8 (ganeshie8):

you're assuming the infinite geometric series converges

OpenStudy (anonymous):

This is the tricky thing about infinite series: normal rules of algebra and math in general don't properly apply.

OpenStudy (anonymous):

@ganeshie8 I read "How not to be wrong" , this is one of insane problem

ganeshie8 (ganeshie8):

however -1 is not a nonsense number :) read below : http://en.wikipedia.org/wiki/Ramanujan_summation

OpenStudy (anonymous):

Yeah!! It's so nonsense, right? sum of a bunch of positive number = a negative one!! ha!!

ganeshie8 (ganeshie8):

your work is equivalent to using the infinite *converging* series formula : \[1+2+4+8 + 16 + \cdots = \dfrac{1}{1-2} = - 1\]

ganeshie8 (ganeshie8):

its mostly nonsense, but ramanujan assigned some numbers to these diverging series and analyzed them... they're fun to read.... go through that wiki link :)

OpenStudy (anonymous):

Thank you. This book makes me craaaazy

ganeshie8 (ganeshie8):

"How not to be wrong" is the book name is it ? looks like an interesting read xD

OpenStudy (anonymous):

will look for that in library thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

is it given why is it not wrong in the book

OpenStudy (ikram002p):

hehe the sum of all even numbers converge to negative result . also there is many approches like this i get to know xD amazing i guess

ganeshie8 (ganeshie8):

\[\large 1-1+1+1-1+\cdots = \dfrac{1}{1-(-1)} = \dfrac{1}{2}\]

OpenStudy (ikram002p):

wowi read this yesturday O.O 1-1+1-1+1-1+...

ganeshie8 (ganeshie8):

\(\large S = 1-1+1-1+\cdots \) \(\large S-1 = -1+1-1+\cdots \) \(\large ~~~~~~~~ = -(1-1+1-\cdots) \) \(\large ~~~~~~~~ = -S \) \(\large \implies 2S =1 \) \(\large \implies S =\dfrac{1}{2} \)

OpenStudy (ikram002p):

but i read something diffrente it was to convice that 1-1+1-1+1-1+1_.. is diverge

ganeshie8 (ganeshie8):

ofcourse it diverges and we can prove it using alternating series test

ganeshie8 (ganeshie8):

but all these calculations are for analyzing diverging series only as we don't want to forget them by simply saying they diverge.

OpenStudy (ikram002p):

ik ^^ i only said what was the topic i read it in xD